OpenStudy (anonymous):

In simplest form, what is the ratio of triangles to circles? A set of shapes is arranged as follows: triangle, circle, square, circle, circle, square, triangle, circle.

OpenStudy (carlyleukhardt):

1/3

OpenStudy (carlyleukhardt):

you have 2:4 2 triangles and 4 circles so total objects jsut counting triangles and circles = 2+4 = 6 so triangles out of total objects = 2/6 = 1/3